OpenStudy (butterflydreamer):

Did i do this correctly? (working out below) :) Find the general solution of A_n=1.0105A_(n-1).

OpenStudy (butterflydreamer):

\[A_n = 1.0105A _{n-1}\] The first step would be to move everything so it equals to 0 right? \[A_n - 1.0105A_{n-1} = 0\] And then i'm not too sure how to continue from here

OpenStudy (butterflydreamer):

Do we just assume that the solution is of the form \[x_n = Cm^n\] So we sub it into the equation: \[A_n - 1.0105A_{n-1} = 0\] \[Cm^n - 1.0105Cm^{n-1} = 0\] Then divide through by Cm^n-1 ?? So, \[m-1.0105 = 0 \rightarrow m = 1.0105\] then sub m into \[x_n = Cm^n\] Therefore, \[A_n = C(1.0105)^n\] is the general solution?

OpenStudy (anonymous):

yes it is , you are correct

OpenStudy (butterflydreamer):

excellent, thank you :)!